# Algebraic power series and their automatic complexity II: modulo prime powers

### v2

@@ 8. Orbit size under $\lambda_{0,0}$
and this tends to 00 as $p$, $\alpha$, $h$, or $d$ tends to infinity and the others remain constant. ∎
@@ 9. Diagonals of rational functions
In this section, we widen our scope from algebraic series to diagonals of multivariate rational functions. Over a field of nonzero characteristic, Furstenberg [10] showed that the diagonal of a multivariate rational function is algebraic. This is a special case of the following result due to Denef and Lipshitz [8, Theorem 6.2 and Remark 6.6].
@@
$p^{N}+p^{N-\alpha((\alpha+1)(h+d)-2)/2}\mathcal{L}(h,h,d,d)\\ +\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{b}},u_{\textnormal{t}})+\left\lceil\log_{p}\max(h,d)\right\rceil+\max(\alpha,2(\alpha-1))+\tfrac{p^{u}-1}{p-1}.$
@@
$\mathcal{W}$ :=$\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in W_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}$
$\mathcal{V}$ :=$\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in V_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}.$
@@
10. Compatibility of automata
Finally, we return to the setting of Theorems 1 and 2, where $F$ is the Furstenberg series associated with a polynomial $P$. When we vary $\alpha$, the automata $\mathcal{M}_{p^{\alpha}}$ support an inverse limit, as described by the authors [18]. In this section, we show that the base-$\frac{p}{Q}$ representations of states in $\mathcal{M}_{p^{\alpha+1}}$ project onto those of states in $\mathcal{M}_{p^{\alpha}}$. This gives a new proof of the existence of the inverse limit, with explicit descriptions of the states as sequences of base-$\frac{p}{Q}$ digits. We begin with an example.
Example 52.
Notation.
Theorem 53.
Proof.
For $m=0$, we have $T_{0}^{\prime}=U_{0,0}\equiv\Lambda_{r,0}\!\left(T_{k}\hat{Q}^{-1}\right)\hat{Q}\mod p$. Therefore the 00th digits satisfy $\operatorname{dig}_{0}(\lambda_{r,0}^{(\alpha)}(S^{(\alpha)}))=\operatorname{dig}_{0}(\lambda_{r,0}^{(\beta)}(S^{(\beta)}))$.
Therefore
Corollary 54.
Proof.
[…]
It remains to show that for each $k\in\{0,1,\dots,\beta-1\}$, we have $\operatorname{dig}_{k}(S_{0}^{(\alpha)})=\operatorname{dig}_{k}(S_{0}^{(\beta)})$. From the proof of Theorem 17, the 00th digit of $S_{0}^{(\alpha)}$ is $T_{0}=\left(y\frac{\partial P}{\partial y}\bmod p\right)$, and this is independent of $\alpha$. Recursively, for $k\geq 1$, the $k$th digit of $S_{0}^{(\alpha)}$ is
@@ Acknowledgments
We thank Frits Beukers and Armin Straub for discussions about their related work.
@@
Table 2 contains data on maximal orbit sizes under $\lambda_{0,0}$. For these searches, we assume that the coefficients in $P$ belong to $\{0,1,\dots,p-1\}$ to make the search space more accessible. In practice, this restriction does not seem to be consequential. For $p^{\alpha}\in\{4,8\}$, $d=1$, and $h\leq 4$, only one polynomial with a coefficient outside this range results in a larger orbit size, namely $(2x+1)y+x$, which yields orbit size $3$ modulo $4$ and also $3$ modulo $8$. For $p^{\alpha}\in\{4,8\}$, $d=2$, and $h\leq 2$, no polynomials result in a larger orbit size.
Surprisingly, all but one of the polynomials that maximize the orbit size for $p^{\alpha}=4$ in Table 2 also maximize the orbit size for $p^{\alpha}=8$. These polynomials are therefore good candidates for obtaining maximal orbit sizes for larger values of $\alpha$. For example, let $P=x^{2}y^{2}+(x^{2}+x+1)y+x^{2}$. For each $\alpha\in\{1,2,\dots,14\}$, the transient length under $\lambda_{0,0}$ is $\alpha+1$ and the eventual period length is $2^{\alpha+1}$, leading one to conjecture that the orbit size is exactly $12^{\alpha+1}+\alpha+$ 1 for all $\alpha\geq 1$. This perhaps suggests the true growth rate, indicating a possible direction for future work.
@@
$\begin{array}[]{c|c|c|c|c}h&d&P&\text{orbit size}&\text{bound}\\ \hline\cr\hline\cr 1&1&y+x&2&9\\ 2&1&y+x^{2}+x&4&135\\ 3&1&y+x^{3}+x^{2}&4&3080\\ 4&1&y+x^{4}+x&5&65545\\ 5&1&(x^{5}+x+1)y+x&7&1572874\\ 6&1&(x^{5}+x+1)y+x^{6}+x&8&25165834\\ 7&1&(x^{7}+x^{6}+x^{2}+x+1)y+x&13&805306378\\ 8&1&(x^{8}+x^{3}+1)y+x&16&1.61\times 10^{10}\\ 9&1&(x^{9}+x^{2}+1)y+x&21&3.43\times 10^{11}\\ 10&1&(x^{10}+x^{6}+x^{3}+x^{2}+1)y+x&31&8.24\times 10^{12}\\ \hline\cr 1&2&xy^{2}+(x+1)y+x&4&133\\ 2&2&x^{2}y^{2}+(x^{2}+x+1)y+x^{2}&11&2097159\\ 3&2&(x^{3}+x+1)y^{2}+(x^{3}+1)y+x^{3}&25&1.03\times 10^{11}\\ 4&2&(x^{4}+x^{2}+x)y^{2}+(x^{4}+x+1)y+x^{4}&50&1.68\times 10^{15}\\ 5&2&(x^{5}+x^{3}+1)y^{2}+(x^{5}+x+1)y+x&121&4.61\times 10^{19}\\ 6&2&(x^{6}+x^{4}+x)y^{2}+(x^{5}+x+1)y+x^{3}&122&7.55\times 10^{23}\\ 7&2&(x^{7}+x+1)y^{2}+(x^{7}+x^{6}+x^{5}+x+1)y+x&337&1.73\times 10^{28}\end{array}$
@@
$\begin{array}[]{c|c|c|c|c}h&d&P&\text{orbit size}&\text{bound}\\ \hline\cr\hline\cr 1&1&y+x&2&4104\\ 2&1&y+x^{2}+x&5&1.37\times 10^{11}\\ 3&1&y+x^{3}+x^{2}&5&3.45\times 10^{18}\\ 4&1&y+x^{4}+x&6&7.73\times 10^{25}\\ 5&1&(x^{5}+x+1)y+x&8&1.94\times 10^{33}\\ 6&1&(x^{5}+x+1)y+x^{6}&9&3.26\times 10^{40}\\ 7&1&(x^{7}+x^{6}+x^{2}+x+1)y+x&14&1.09\times 10^{48}\\ 8&1&(x^{8}+x^{3}+1)y+x&17&2.29\times 10^{55}\\ 9&1&(x^{9}+x^{2}+1)y+x&22&5.14\times 10^{62}\\ 10&1&(x^{10}+x^{6}+x^{3}+x^{2}+1)y+x&32&1.29\times 10^{70}\\ \hline\cr 1&2&xy^{2}+(x+1)y+x&5&1.37\times 10^{11}\\ 2&2&x^{2}y^{2}+(x^{2}+x+1)y+x^{2}&20&1.01\times 10^{31}\\ 3&2&(x^{3}+x+1)y^{2}+(x^{3}+1)y+x^{3}&50&2.24\times 10^{51}\\ 4&2&(x^{4}+x^{2}+x)y^{2}+(x^{4}+x+1)y+x^{4}&99&1.65\times 10^{71}\\ 5&2&(x^{5}+x^{3}+1)y^{2}+(x^{5}+x+1)y+x&242&2.03\times 10^{91}\\ 6&2&(x^{6}+x^{4}+x)y^{2}+(x^{5}+x+1)y+x^{3}&243&1.50\times 10^{111}\\ 7&2&(x^{7}+x+1)y^{2}+(x^{7}+x^{6}+x^{5}+x+1)y+x&674&1.55\times 10^{131}\end{array}$
@@
$\begin{array}[]{c|c|c|c|c}h&d&P&\text{orbit size}&\text{bound}\\ \hline\cr\hline\cr 1&1&y+x&2&14\\ 2&1&y+x^{2}+x&3&1464\\ 3&1&y+x^{3}+x&4&177154\\ 4&1&(2x^{4}+x+1)y+x&5&19131884\\ 5&1&(x^{5}+2x^{2}+1)y+x&7&2324522942\\ 6&1&(x^{5}+2x^{2}+1)y+x^{6}&8&1.88\times 10^{11}\\ \hline\cr 1&2&xy^{2}+y+x&4&1463\\ 2&2&(x^{2}+x+1)y^{2}+y+x^{2}+x&20&6973568808\\ 3&2&(x^{3}+x^{2}+x+2)y^{2}+(x+1)y+x^{3}+x&57&1.00\times 10^{17}\\ 4&2&(x^{3}+2x+1)y^{2}+(x^{4}+1)y+x&218&4.78\times 10^{23}\end{array}$
Table 2. Polynomials with coefficients in $\{0,1,\dots,p-1\}$ for which the initial state achieves the maximum orbit size under $\lambda_{0,0}$ for given values of $p$, $\alpha$, $h$, and $d$. The final column contains the value of $p^{N-\alpha(\alpha+1)(h+d-1)/2}\mathcal{L}(h,d,d)+\left\lfloor\log_{p}\max\!\left(p^{\alpha-1}h,p^{\alpha-1}(d-1)\right)\right\rfloor+1+\left\lceil\log_{p}\max(h,d-1)\right\rceil+\max(\alpha,2(\alpha-1))+1$ from Theorem 2.

### v3

@@ 10. Non-Furstenberg series
and this is in $o(1)p^{M}$ as any of $p$, $\alpha$, $h^{\prime}$, or $d$ tends to infinity and the others remain constant. Therefore $\lvert{\operatorname{orb}_{\Lambda_{0}}(G\bmod p^{\alpha})}\rvert$ is in $o(1)p^{N}$ as any of $p$, $\alpha$, $h$, or $d$ tends to infinity and the others remain constant.
It remains to bound $\lvert{\operatorname{orb}_{\Lambda_{p-1}}(G_{q}\bmod p^{\alpha})}\rvert$ for each $q\in\{1,2,\dots,e\}$, where $G_{q}$ is defined in Case 1 in the proof of Lemma 53. If $e=0$, then there are no such terms and we are done. Assume $e\geq 1$. Since $G$ is a Furstenberg series, Theorem 4 implies that $G$ is the diagonal of a rational function with denominator $Q_{e}\mathrel{\mathop{:}}=P_{e}/(x^{s}z)$. Furthermore, as in Section 2, this implies that every element of $\ker_{p}((b(n)\bmod p^{\alpha})_{n\geq 0})$ is the coefficient sequence of the diagonal of a rational function with denominator $Q_{e}^{p^{\alpha-1}}$. In particular, $G_{q}$ is the diagonal of such a rational function; let its numerator be $S_{q}$. Therefore we can access each $\Lambda_{p-1}(G_{q})$ by working with the following operator. Define $\lambda_{p-1,0}\colon\mathcal{R}_{p^{\alpha}}[x,z,z^{-1}]\to\mathcal{R}_{p^{\alpha}}[x,z,z^{-1}]$ by
$\lambda_{p-1,0}(S)\mathrel{\mathop{:}}=\Lambda_{p-1,0}\!\left(SQ_{e}^{p^{\alpha}-p^{\alpha-1}}\right).$
Since by assumption $\deg_{x}P_{e}=\deg_{x}(P_{e}\bmod p)$ and $\deg_{z}P_{e}=\deg_{z}(P_{e}\bmod p)$, the value of $u$ in Proposition 26 is $1$. Furthermore, inspecting Case 1 in the proof of Lemma 53, we see that we apply at least one Cartier operator $\Lambda_{p-1}$ to each of the series $G_{q}\bmod p^{\alpha}$ defined there. By Corollary 24, the base-$\frac{p}{Q_{e}}$ representation of $S_{q}Q_{e}^{p^{\alpha}-p^{\alpha-1}}$ belongs to the space $\mathcal{W}$ defined by
$\mathcal{W}\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in W_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}$
where
$W_{k}\mathrel{\mathop{:}}=\left\{\sum_{i=0}^{(k+1)h^{\prime}-1}\sum_{j=\max(-k,-i)}^{(k+1)(d-1)}c_{i,j}x^{i}z^{j}:\text{$c_{i,j}\in D$ for each $i,j$}\right\}$
analogously to Equation (8). Accordingly, Equation (5) implies that the size of the basis of $\mathcal{W}$ is $M$.
We modify the part of the proof in Theorem 2 where we bounded an orbit size under $\lambda_{0,0}$ to bound an orbit size under $\lambda_{p-1,0}$. The key idea there was to use the results in Section 6 in order to restrict the spaces in which the base-$\frac{p}{Q}$ representations of elements in the orbit live. We moved from the larger space $\mathcal{V}$ defined in Equation (9) to the smaller space $\mathcal{V}^{\circ}$ defined in Equation (22). Here we do something similar; namely, we move from $\mathcal{W}$ to a space $\mathcal{W}^{\circ}$ defined by
$\mathcal{W}^{\circ}\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in W_{k}^{\circ}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}$
where
$W_{k}^{\circ}\mathrel{\mathop{:}}=\left\{\sum_{i=0}^{(k+1)h^{\prime}-1}\sum_{j=\max(-k,-i)}^{(k+1)(d-1)-1}c_{i,j}x^{i}z^{j}:\text{$c_{i,j}\in D$ for each $i,j$}\right\}$
for each $k\in\{0,1,\dots,\alpha-1\}$. (Since only some of the results in Section 6 extend to $\lambda_{p-1,0}$, the space $\mathcal{W}^{\circ}$ is slightly larger than $\mathcal{V}^{\circ}$ as defined in Equation (22).) We have
$\dim\mathcal{W}^{\circ}$ $=\sum_{k=0}^{\alpha-1}\left((k+1)h^{\prime}\cdot(k+1)(d-1)+\sum_{j=-k}^{-1}\left((k+1)h^{\prime}+j\right)\right)$
$=\tfrac{1}{6}\alpha(\alpha+1)((2h^{\prime}d-1)\alpha+h^{\prime}d-3h^{\prime}+1)$
$=M-\tfrac{1}{2}\alpha(\alpha+1)h^{\prime}.$
The results that allow us to move from $\mathcal{W}$ to the smaller space $\mathcal{W}^{\circ}$ are the following. The first result is a variant of the third part of Proposition 29. Using the notation from Section 6, let $R=\pi_{z,d-1}(Q_{e})$. Then for all $S\in\mathcal{R}_{p^{\alpha}}[x,z,z^{-1}]$ with degree at most $p^{\alpha-1}(d-1)$, we have
$\pi_{z,p^{\alpha-1}(d-1)}(\lambda_{p-1,0}(S))=\lambda_{p-1}(\pi_{z,p^{\alpha-1}(d-1)}(S)).$
The second result is a variant of the third part of Theorem 30. Namely, if $S$ is a state in $\operatorname{val}_{p/Q_{e}}(\mathcal{V})$ and $R=\pi_{z,d-1}(Q_{e})$, then $\operatorname{pr}_{z,d-1}(\operatorname{rep}_{p/Q_{e}}(S))=\operatorname{rep}_{p/R}(\pi_{z,p^{\alpha-1}(d-1)}(S))$. With these two results, the proof in Theorem 2 also applies to $\lambda_{p-1,0}$, giving
$\lvert{\operatorname{orb}_{\lambda_{p-1,0}}(S_{q})}\rvert\leq p^{M-\alpha(\alpha+1)h^{\prime}/2}g(h^{\prime})+\text{small terms}.$
We now proceed as in the proof of Theorem 1 to compare the size of $\operatorname{orb}_{\lambda_{p-1,0}}(S_{q})$ to $M$. Ignoring the small terms, we have
$\frac{\lvert{\operatorname{orb}_{\lambda_{p-1,0}}(S_{q})}\rvert}{p^{M}}\leq\frac{p^{M-\alpha(\alpha+1)h^{\prime}/2}g(h^{\prime})}{p^{M}}=\frac{g(h^{\prime})}{p^{\alpha(\alpha+1)h^{\prime}/2}}=\frac{e^{(1+\epsilon(h^{\prime}))\sqrt{h^{\prime}\log h^{\prime}}}}{p^{\alpha(\alpha+1)h^{\prime}/2}}.$
This expression tends to 0 as $p$, $\alpha$, or $h^{\prime}$ tends to infinity and the others remain constant, so $\lvert{\operatorname{orb}_{\lambda_{p-1,0}}(S_{q})}\rvert\in o(1)p^{M}$. Since $e\geq 1$, if $h$ or $d$ tends to infinity then $h^{\prime}$ tends to infinity. Since $M\leq N$, then $\lvert{\operatorname{orb}_{\lambda_{p-1,0}}(S_{q})}\rvert\in o(1)p^{N}$ as $p$, $\alpha$, $h$, or $d$ tend to infinity and the others remain constant. ∎
@@ 11. Diagonals of rational functions
In this section, we widen our scope from algebraic series to diagonals of multivariate rational functions. Over a field of nonzero characteristic, Furstenberg [12] showed that the diagonal of a multivariate rational function is algebraic. This is a special case of the following result due to Denef and Lipshitz [10, Theorem 6.2 and Remark 6.6].
@@
$p^{N}+p^{N-\alpha((\alpha+1)(h+d)-2)/2}\mathcal{L}(h,h,d,d)\\ +\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{b}},u_{\textnormal{t}})+\left\lfloor\log_{p}\max(h,d)\right\rfloor+2\alpha-1+\tfrac{p^{u}-1}{p-1}.$
@@
Using $\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{b}},u_{\textnormal{t}})\leq\left\lfloor\log_{p}\max(h,d)\right\rfloor+\alpha$, we obtain the simpler bound
$p^{N}+p^{N-\alpha((\alpha+1)(h+d)-2)/2}\mathcal{L}(h,h,d,d)+2\left\lfloor\log_{p}\max(h,d)\right\rfloor+3\alpha-1+\tfrac{p^{u}-1}{p-1}.$
@@
$\mathcal{W}$ $\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in W_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}$
@@ Acknowledgments
We thank Frits Beukers and Armin Straub for discussions about their related work. We also thank the reviewers for constructive feedback.
@@
Table 2 contains data on maximal orbit sizes under $\lambda_{0,0}$. For these searches, we assume that the coefficients in $P$ belong to $\{0,1,\dots,p-1\}$ to make the search space more accessible. In practice, this restriction does not seem to be consequential. For $p^{\alpha}\in\{4,8\}$, $d=1$, and $h\leq 4$, only one polynomial with a coefficient outside this range results in a larger orbit size, namely $(2x+1)y+x$, which yields orbit size $3$ modulo 44 and also $3$ modulo $8$. For $p^{\alpha}\in\{4,8\}$, $d=2$, and $h\leq 2$, no polynomials result in a larger orbit size.
Surprisingly, all but one of the polynomials that maximize the orbit size for $p^{\alpha}=4$ in Table 2 also maximize the orbit size for $p^{\alpha}=8$. These polynomials are therefore good candidates for obtaining maximal orbit sizes for larger values of $\alpha$. For example, let $P=x^{2}y^{2}+(x^{2}+x+1)y+x^{2}$. For each $\alpha\in\{1,2,\dots,14\}$, the transient length under $\lambda_{0,0}$ is $\alpha+1$ and the eventual period length is $2^{\alpha+1}$, leading one to conjecture that the orbit size is exactly $2^{\alpha+1}+\alpha+1$ for all $\alpha\geq 1$. This perhaps suggests the true growth rate, indicating a possible direction for future work.
@@
$\begin{array}[]{c|c|c|c|c}h&d&P&\text{orbit size}&\text{bound}\\ \hline\cr\hline\cr 1&1&y+x&2&10\\ 2&1&y+x^{2}+x&4&136\\ 3&1&y+x^{3}+x^{2}&4&3080\\ 4&1&y+x^{4}+x&5&65546\\ 5&1&(x^{5}+x+1)y+x&7&1572874\\ 6&1&(x^{5}+x+1)y+x^{6}+x&8&25165834\\ 7&1&(x^{7}+x^{6}+x^{2}+x+1)y+x&13&805306378\\ 8&1&(x^{8}+x^{3}+1)y+x&16&1.61\times 10^{10}\\ 9&1&(x^{9}+x^{2}+1)y+x&21&3.43\times 10^{11}\\ 10&1&(x^{10}+x^{6}+x^{3}+x^{2}+1)y+x&31&8.24\times 10^{12}\\ \hline\cr 1&2&xy^{2}+(x+1)y+x&4&135\\ 2&2&x^{2}y^{2}+(x^{2}+x+1)y+x^{2}&11&2097160\\ 3&2&(x^{3}+x+1)y^{2}+(x^{3}+1)y+x^{3}&25&1.03\times 10^{11}\\ 4&2&(x^{4}+x^{2}+x)y^{2}+(x^{4}+x+1)y+x^{4}&50&1.68\times 10^{15}\\ 5&2&(x^{5}+x^{3}+1)y^{2}+(x^{5}+x+1)y+x&121&4.61\times 10^{19}\\ 6&2&(x^{6}+x^{4}+x)y^{2}+(x^{5}+x+1)y+x^{3}&122&7.55\times 10^{23}\\ 7&2&(x^{7}+x+1)y^{2}+(x^{7}+x^{6}+x^{5}+x+1)y+x&337&1.73\times 10^{28}\end{array}$
@@
$\begin{array}[]{c|c|c|c|c}h&d&P&\text{orbit size}&\text{bound}\\ \hline\cr\hline\cr 1&1&y+x&2&4105\\ 2&1&y+x^{2}+x&5&1.37\times 10^{11}\\ 3&1&y+x^{3}+x^{2}&5&3.45\times 10^{18}\\ 4&1&y+x^{4}+x&6&7.73\times 10^{25}\\ 5&1&(x^{5}+x+1)y+x&8&1.94\times 10^{33}\\ 6&1&(x^{5}+x+1)y+x^{6}&9&3.26\times 10^{40}\\ 7&1&(x^{7}+x^{6}+x^{2}+x+1)y+x&14&1.09\times 10^{48}\\ 8&1&(x^{8}+x^{3}+1)y+x&17&2.29\times 10^{55}\\ 9&1&(x^{9}+x^{2}+1)y+x&22&5.14\times 10^{62}\\ 10&1&(x^{10}+x^{6}+x^{3}+x^{2}+1)y+x&32&1.29\times 10^{70}\\ \hline\cr 1&2&xy^{2}+(x+1)y+x&5&1.37\times 10^{11}\\ 2&2&x^{2}y^{2}+(x^{2}+x+1)y+x^{2}&20&1.01\times 10^{31}\\ 3&2&(x^{3}+x+1)y^{2}+(x^{3}+1)y+x^{3}&50&2.24\times 10^{51}\\ 4&2&(x^{4}+x^{2}+x)y^{2}+(x^{4}+x+1)y+x^{4}&99&1.65\times 10^{71}\\ 5&2&(x^{5}+x^{3}+1)y^{2}+(x^{5}+x+1)y+x&242&2.03\times 10^{91}\\ 6&2&(x^{6}+x^{4}+x)y^{2}+(x^{5}+x+1)y+x^{3}&243&1.50\times 10^{111}\\ 7&2&(x^{7}+x+1)y^{2}+(x^{7}+x^{6}+x^{5}+x+1)y+x&674&1.55\times 10^{131}\end{array}$
@@
$\begin{array}[]{c|c|c|c|c}h&d&P&\text{orbit size}&\text{bound}\\ \hline\cr\hline\cr 1&1&y+x&2&15\\ 2&1&y+x^{2}+x&3&1464\\ 3&1&y+x^{3}+x&4&177155\\ 4&1&(2x^{4}+x+1)y+x&5&19131884\\ 5&1&(x^{5}+2x^{2}+1)y+x&7&2324522942\\ 6&1&(x^{5}+2x^{2}+1)y+x^{6}&8&1.88\times 10^{11}\\ \hline\cr 1&2&xy^{2}+y+x&4&1464\\ 2&2&(x^{2}+x+1)y^{2}+y+x^{2}+x&20&6973568808\\ 3&2&(x^{3}+x^{2}+x+2)y^{2}+(x+1)y+x^{3}+x&57&1.00\times 10^{17}\\ 4&2&(x^{3}+2x+1)y^{2}+(x^{4}+1)y+x&218&4.78\times 10^{23}\end{array}$
Table 2. Polynomials with coefficients in $\{0,1,\dots,p-1\}$ for which the initial state achieves the maximum orbit size under $\lambda_{0,0}$ for given values of $p$, $\alpha$, $h$, and $d$. The final column contains the value of $p^{N-\alpha(\alpha+1)(h+d-1)/2}\mathcal{L}(h,d,d)+\left\lfloor\log_{p}\max\!\left(p^{\alpha-1}h,p^{\alpha-1}(d-1)\right)\right\rfloor+1+\left\lfloor\log_{p}\max(h,d)\right\rfloor+2\alpha-1+1$ from Theorem 2.